# AN ACCELERATED PRIMAL DUAL ALGORITHM WITH BACKTRACKING FOR DECENTRALIZED CONSTRAINED OPTIMIZATION

### v1

@@ 8. Conclusion
In this work, we consider cooperative multi-agent constrained consensus optimization problems over an undirected network of agents. Agents have local objective functions with a composite structure—each is the sum of smooth and nonsmooth terms—and each agent is subject to its own private convex conic constraints. We consider the setting when only agents connected by an edge can directly communicate to exchange large-volume data vectors via a high-speed, short-range protocol (e.g., WiFi), and when the network supports one-hop, low-rate information exchange beyond immediate neighbors, as in the LoRaWAN protocol. In this setting, unlike existing decentralized primal–dual methods that require knowledge of Lipschitz constants, we propose a new algorithm, D-APDB, which adapts to local smoothness via a distributed backtracking step-size search. We obtain $\mathcal{O}(1/K)$ convergence rate guarantees for suboptimality, infeasiability and consensus violation. When nodes have private constraints for which projections are expensive to compute, D-APDB is, to the best of our knowledge, the first distributed method with backtracking that attains the optimal convergence rate for constrained composite convex optimization problems. We also provide a variant of our algorithm, D-APDB0, for problems that are either unconstrained or with simple constraint sets for which projections are cheap to compute. D-APDB0 can also achieve the same $\mathcal{O}(1/K)$ rate for suboptimality and consensus violation. Finally, we present numerical experiments demonstrating the performance of D-APDB0 and D-APDB on distributed $\ell_{1}$-norm-regularized QP and QCQP problems.

### v2

@@ 8. Conclusion
In this work, we consider cooperative multi-agent constrained consensus optimization problems over an undirected network of agents. Agents have local objective functions with a composite structure—each is the sum of smooth and nonsmooth terms—and each agent is subject to its own private convex conic constraints. We consider the setting when only agents connected by an edge can directly communicate to exchange large-volume data vectors via a high-speed, short-range protocol (e.g., WiFi), and when the network supports one-hop, low-rate information exchange beyond immediate neighbors, as in the LoRaWAN protocol. In this setting, unlike existing decentralized primal–dual methods that require knowledge of Lipschitz constants, we propose a new algorithm, D-APDB, which adapts to local smoothness via a distributed backtracking step-size search. We obtain $\mathcal{O}(1/K)$ convergence rate guarantees for suboptimality, infeasibility and consensus violation. When nodes have private constraints for which projections are expensive to compute, D-APDB is, to the best of our knowledge, the first distributed method with backtracking that attains the optimal convergence rate for constrained composite convex optimization problems. We also provide a variant of our algorithm, D-APDB0, for problems that are either unconstrained or with simple constraint sets for which projections are cheap to compute. D-APDB0 can also achieve the same $\mathcal{O}(1/K)$ rate for suboptimality and consensus violation. Finally, we present numerical experiments demonstrating the performance of D-APDB0 and D-APDB on distributed $\ell_{1}$-norm-regularized QP and QCQP problems, and distributed primal SVM training.
Acknowledgements
This paper builds on momentum acceleration, and we dedicate it to Prof. Yurii Nesterov (UCLouvain, Belgium), whose seminal contributions to the development and analysis of optimization algorithms have been profoundly influential in this field. This research was supported in part by the Office of Naval Research under award number N00014-24-1-2628 and N00014-24-1-2666.